A firm has $40 million in sales, a lerner index of .45, and a marginal cost of $25, and competes against 200 other firms inits relevant market. a.) What price does this firm charge its customers? b.) By what factor does this firm markup it price or marginal cost?
Answer #1

Answer

(a)

Formula :

Lerner Index = (P - MC)/P where P = Price and MC = Marginal Cost

Here, it is given that MC = Marginal cost = 25, Lerner Index = 0.45

=> Lerner Index = (P - MC)/P --------------------(1)

=> 0.45 = (P - 25)/P

=> 0.45P = P - 25

=> 0.55P = 25

=> P = 45.45.

Hence, this firm will charge its customers $45.45(approx)

(b)

From (1) we have Lerner Index = (P - MC)/P and it is given that Lerner index = 0.45

=> 0.45 = (P - MC)/P

=> 0.45P = P - MC

=> 0.55P = MC

=> P = 1.818*MC.

Hence, factor by which this firm markup it price or marginal cost = 1.818